OpenStudy (anonymous):

A computer screen has a width of 15 inches and a height of 11 inches. Approximately what is the length of the screen’s diagonal?

OpenStudy (anonymous):

L*W=D hope this helps

OpenStudy (anonymous):

diagonal, use pythagoras

OpenStudy (anonymous):



OpenStudy (anonymous):

\[11^2+15^2=d^2\] so \[d=\sqrt{15^2+11^2}\] whatever that is

OpenStudy (anonymous):

a squared*b Squared =c squared. correct?

OpenStudy (anonymous):

plus, not times

OpenStudy (anonymous):

oh ok